OpenStudy (anonymous):

prove that pressure is equal to height*density*9.8

OpenStudy (anonymous):

height matters because of how much mass is on top of you, density is the density of molecules which determines their weight and 9.8 is the gravitational force which is obvious

OpenStudy (anonymous):

The pain you'd feel is the direct result of the force that brought about ... I get that the height of the beaker is H = 1.36 meters and the height of .... Let m (or mu) be the linear mass density of the chain. ... In the end all that matters is the weight of the chain, and the force at the top. ..... v_w=4.0, 4.7, 7.5, 9.8 (m/s),.

OpenStudy (anonymous):

actually the pressure is Density*gravity*height as \[f=mass \times acceleration\] \[mass= density \times area \times height\] \[f=pressure \times area\] area gets cancelled and pressure=density*g*h

OpenStudy (anonymous):

well i tried this. Pressure= Height*Density*9.8 is the same as Pressure= Height*Density*G So, now i do math's with Dimensions. [M][L]^-1[T]^-2= [L][M][L]^-3[L][T]^-2 By simplifying the [L]'s from the right side i'd get: [M][L]^-1[T]^-2=[M][L]^-1[T]^-2 Not sure if it's the end. But so far it's 100% Mathematically correct!!! :)

OpenStudy (anonymous):

the preuve is : That the lord of pressure says that P = F / S and by knowing that density D = M / V so D * H = M / S also that the force F = M * A we can concluse other equation for the pressure and that is P = H * D * 9.8 .